OpenStudy (anonymous):

3 times the measure of an angle is 14 less than the measure of its complement. What is the measure of the angle? 19° 57° 71° 76°

OpenStudy (raden):

let that angle is x, therefore its complement be 90 - x 3x = 90 - x - 14 simplify then solve for x

OpenStudy (anonymous):

3x=(90-x)-14 solve for x